OpenStudy (anonymous):

A wooden fence 4 feet high and 200 feet long is to be painted. How many gallons of paint are needed if one gallon covers 300 square feet? Assume that the paint is not sold in partial quantities.

OpenStudy (anonymous):

@One098

OpenStudy (anonymous):

@Nurali need help on this problem i know you are also a mathlete

OpenStudy (anonymous):

hey one098 nurali's problem solveing is at champion he is smart score is 99 cause of his problem solving

OpenStudy (anonymous):

@koadmay

OpenStudy (anonymous):

do you think you can help me on this question

OpenStudy (anonymous):

that helped big time nother meda

OpenStudy (anonymous):

medal

OpenStudy (anonymous):

Find the area of a triangle with base of 10 inches and a height of 16 inches.

OpenStudy (anonymous):

A = 80 in²

OpenStudy (anonymous):

A=80in^2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

ima copy and paste it into text edit and sve it for when i got the work back

OpenStudy (anonymous):

A rhombus contains two pair of parallel sides. If the equation for the line containing one of the sides is y = x - 3, what is the equation of the line opposite line, passing through (0, 4)? y = x - 3 y = x + 4 y = -x + 4

OpenStudy (anonymous):

i dont really know this one

OpenStudy (anonymous):

it is okay neither do i